# YOLO Meets Mixture-of-Experts: Adaptive Expert Routing for Robust Object Detection

### v3

@@ I Introduction
Object detection is one of the most fundamental and challenging problems in computer vision. It aims to identify and localize instances of objects within images or video frames and has received great attention in recent years [zou2023object, zhao2019object, zou2019review].
The introduction of convolutional neural networks (CNNs) revolutionized the field, giving rise to detectors such as R-CNN [RCNN], single-shot detector (SSD) [liu2016ssd], and YOLO [wang2024yolov9, jiang2025odverse33newyoloversion, YoloOG], which achieved remarkable improvements in both accuracy and speed. Among them, YOLOv9 has shown superior performance compared to other models in the YOLO series [jiang2025odverse33newyoloversion].
In parallel with these developments, the Mixture-of-Experts (MoE) [mu2025comprehensivesurveymixtureofexpertsalgorithms] paradigm introduces multiple specialized experts and a routing network that dynamically assigns responsibility for each input. The outputs are combined either softly, as weighted mixtures [shazeer2017], or through hard expert selection [switchTransformers]. In [YoloMoE], a MoE module is integrated into the YOLOv5 [yolov5] backbone to process intermediate feature maps. The MoE framework has also proven effective in tasks such as next-token prediction [switchTransformers] and deepfake detection [negroni2025leveraging]. In many of these applications, the outputs correspond to class probabilities or logits, which can be combined straightforwardly as:
@@
Building on these advances, this paper introduces an MoE architecture based on YOLOv9-T (the tiny version of YOLOv9) [wang2024yolov9], designed to improve detection performance and robustness.
Figure 1: Visualization of MoE within the YOLOv9 architecture, multiple experts process the input image to produce multi-scale feature maps and outputs (class and bounding box logits). Routers at different resolutions (8×8, 16×16, 32×32) generate adaptive routing weights that fuse expert outputs into final detections. The loss is computed between model outputs before Non-maximum suppression (NMS) [Hosang2017LearningNS, wang2024yolov9] and the ground truth, ensuring end-to-end differentiability.
@@ II Proposed MoE-Based YOLOv9 Architecture
In the YOLOv9 framework, each input image is processed through $I=3$ feature maps $F_{i}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}$ corresponding to spatial resolutions $i\in\{8,16,32\}$, where $B$ denotes the batch size, $C_{i}$ the number of channels, and $(H_{i},W_{i})$ the spatial dimensions of the $i$-th level. Each map is responsible for detecting objects at a corresponding scale, and instead of directly regressing the bounding box coordinates, the model predicts the distances from each cell center to the four sides of the bounding box (left, top, bottom, right). These distances are discretized into NN bins (typically 16) and treated as classification targets, with coordinates obtained by the expectation over predicted probabilities. This discrete formulation enables consistent probabilistic fusion across experts (YOLOv9 models).
In this framework, each feature level produces a prediction map containing both bounding box and class logits. Specifically, the map outputs $y_{\text{box}}\in\mathbb{R}^{4\times N}$ for bounding box regression and $y_{\text{cls}}\in\mathbb{R}^{n_{c}}$ for object classification, where NN denotes the number of discrete distance classes used in bounding box regression, and $n_{c}$ denotes the number of object classes.
Our design follows the Enhanced MoE method proposed in [negroni2025leveraging], where the routing decision is based on features extracted by individual experts rather than on the raw input image. Building on this multi-scale structure of YOLOv9, the resulting feature representations at each level serve as inputs to the MoE routing mechanism.
@@
To prevent the router from collapsing onto a single expert, a load balancing loss inspired by [switchTransformers] is incorporated. The total loss is defined as:
@@ III Experiments and results
We conducted experiments on the COCO [cocoDataset] and VisDrone [Zhu2020DetectionAT] datasets, which represent everyday scenes and aerial imagery, respectively, focusing on the Person, Vehicle, Bicycle, and Motorcycle classes (the classes shared by both datasets). The models were trained for 5050 epochs using the default hyperparameters of the YOLOv9 framework [wang2024yolov9], and our architecture employed two experts based on YOLOv9-T, with $\lambda_{\text{lb}}=0.5$. As in [negroni2025leveraging], the experts were initialized from weights pretrained on their datasets (COCO or VisDrone), enabling dataset-specific specialization, while the routing modules were trained from scratch. The best model was selected according to the mean Average Precision (mAP) metric. The evaluation was conducted using mAP and Average Recall (AR) metrics. As shown in Table I, our model achieved better performance compared to the other YOLOv9-T models on both datasets.

### v4

@@ I Introduction
Object detection is one of the most fundamental and challenging problems in computer vision. It aims to identify and localize instances of objects within images or video frames and has received great attention in recent years [17, 14, 16].
The introduction of convolutional neural networks (CNNs) revolutionized the field, giving rise to detectors such as R-CNN [3], single-shot detector (SSD) [8], and YOLO [13, 6, 11], which achieved remarkable improvements in both accuracy and speed. Among them, YOLOv9 has shown superior performance compared to other models in the YOLO series [6].
In parallel with these developments, the Mixture-of-Experts (MoE) [9] paradigm introduces multiple specialized experts and a routing network that dynamically assigns responsibility for each input. The outputs are combined either softly, as weighted mixtures [12], or through hard expert selection [2]. In [1], a MoE module is integrated into the YOLOv5 [5] backbone to process intermediate feature maps. The MoE framework has also proven effective in tasks such as next-token prediction [2] and deepfake detection [10]. In many of these applications, the outputs correspond to class probabilities or logits, which can be combined straightforwardly as:
@@
Building on these advances, this paper introduces an MoE architecture based on YOLOv9-T (the tiny version of YOLOv9) [13], designed to improve detection performance and robustness.
Figure 1: Visualization of MoE within the YOLOv9 architecture, multiple experts process the input image to produce multi-scale feature maps and outputs (class and bounding box logits). Routers at different resolutions (8×8, 16×16, 32×32) generate adaptive routing weights that fuse expert outputs into final detections. The loss is computed between model outputs before Non-maximum suppression (NMS) [4, 13] and the ground truth, ensuring end-to-end differentiability.
@@ II Proposed MoE-Based YOLOv9 Architecture
In the YOLOv9 framework, each input image is processed through $I=3$ feature maps $F_{i}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}$ corresponding to spatial resolutions $i\in\{8,16,32\}$, where $B$ denotes the batch size, $C_{i}$ the number of channels, and $(H_{i},W_{i})$ the spatial dimensions of the $i$-th level. Each map is responsible for detecting objects at a corresponding scale, and instead of directly regressing the bounding box coordinates, the model predicts the distances from each cell center to the four sides of the bounding box (left, top, bottom, right). These distances are discretized into $N$ bins (typically 16) and treated as classification targets, with coordinates obtained by the expectation over predicted probabilities. This discrete formulation enables consistent probabilistic fusion across experts (YOLOv9 models).
In this framework, each feature level produces a prediction map containing both bounding box and class logits. Specifically, the map outputs $y_{\text{box}}\in\mathbb{R}^{4\times N}$ for bounding box regression and $y_{\text{cls}}\in\mathbb{R}^{n_{c}}$ for object classification, where $N$ denotes the number of discrete distance classes used in bounding box regression, and $n_{c}$ denotes the number of object classes.
Our design follows the Enhanced MoE method proposed in [10], where the routing decision is based on features extracted by individual experts rather than on the raw input image. Building on this multi-scale structure of YOLOv9, the resulting feature representations at each level serve as inputs to the MoE routing mechanism.
@@
To prevent the router from collapsing onto a single expert, a load balancing loss inspired by [2] is incorporated. The total loss is defined as:
@@ III Experiments and results
We conducted experiments on the COCO [7] and VisDrone [15] datasets, which represent everyday scenes and aerial imagery, respectively, focusing on the Person, Vehicle, Bicycle, and Motorcycle classes (the classes shared by both datasets). The models were trained for $50$ epochs using the default hyperparameters of the YOLOv9 framework [13], and our architecture employed two experts based on YOLOv9-T, with $\lambda_{\text{lb}}=0.5$. As in [10], the experts were initialized from weights pretrained on their datasets (COCO or VisDrone), enabling dataset-specific specialization, while the routing modules were trained from scratch. The best model was selected according to the mean Average Precision (mAP) metric. The evaluation was conducted using mAP and Average Recall (AR) metrics. As shown in Table I, our model achieved better performance compared to the other YOLOv9-T models on both datasets.